# G2SDF: Surface Reconstruction from Explicit Gaussians with Implicit SDFs

### v1

@@ 3.3 Scene Representation and Optimization
To address this issue, we back-project the rendered Gaussian depth $\bar{D}$ into 3D to generate dense supervision signals. Further, we also supervise free space to mitigate unexpected noise. More specifically, we first sample $M$ pixels $\{\mathbf{u}=\mathbf{u}_{i}\ |\ \mathbf{u}_{i}\in\mathbb{R}^{2},{\text{for}\ i=1,2,...,M\}}$ from the rendered Gaussian depth map $\bar{D}$. The ray cast from a pixel $\mathbf{u}$ is denoted as $\mathbf{p}(\mathbf{u})=\mathbf{r}_{o}+t\cdot\mathbf{r}_{d}(\mathbf{u})$, where $\mathbf{r}_{o}\in\mathbb{R}^{3}$ is the camera center and $\mathbf{r}_{d}\in\mathbb{R}^{3}$ is the unit direction vector of the ray. For each ray, we sample $K=K_{n}+K_{f}$ points $\{t=t_{k}\ |\ t_{k}\in\mathbb{R},\text{for}\ k=1,2,...,K\}$ from the camera center to the surface, where $K_{n}$ means near surface samples and $K_{f}$ means free space samples. Using Eq. 5, the SDF values of these points can be written as $\bar{s}_{k}$. For near surface samples within a truncation region, i.e. samples where $S_{tr}=\{|\bar{D}(\mathbf{u})-t_{k}|\leq tr\}$, we use the distance between the sampled point $t_{k}$ and its surface $\bar{D}(\mathbf{u}_{j})$ as an approximation of ground truth SDF value for supervision:
@@
For points that are far from the surface $S_{fs}=\{\bar{D}(\mathbf{u})-t_{k}>tr\}$, we apply a free-space loss:
@@
During training, we jointly optimize explicit Gaussians and implicit SDF. The overall loss function for our hybrid scene representation is:
$\mathcal{L}_{scene}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{n}+\lambda_{ns}\mathcal{L}_{ns}+\lambda_{fs}\mathcal{L}_{fs},$ (9)
where we apply the color loss $\mathcal{L}_{c}$ combining L1 with the D-SSIM term, distortion loss $\mathcal{L}_{d}$ and depth-normal loss $\mathcal{L}_{n}$ as defined in GOF [2]. In our approach, Gaussian opacity is directly linked to its position such that we can simplify the 3DGS densification process by removing its opacity reset step and corresponding hyperparameters.
Figure 3: Surface Reconstruction on the DTU Dataset [43]. We show normal maps rendered from the reconstructed meshes. 2DGS [1] produces overly smooth surfaces with floating artifacts and outliers, while GOF [2] struggles to reconstruct flat surfaces and performs poorly in regions with limited view coverage. In contrast, our method generates smooth yet detailed surfaces without floating artifacts, demonstrating its superior ability to represent geometry accurately.
3.4 Geometry Regularization
To ensure high-quality reconstructions on real-world scenes, we further employ an off-the-shelf monocular depth estimator***We use DepthAnything V2 [44] in this work. to obtain dense per-pixel depth priors. To address the scale ambiguity between the estimated depths and the scene, we apply linear regression regularization between rendered Gaussian depth $\bar{D}$ and estimated monocular depth $\hat{D}$:
$\mathcal{L}_{depth}=\sum log(1+|(a*\hat{D}+b)-\bar{D}|).$ (10)
where $a,b$ are learnable parameters for scale and shift. We also compute the gradient of monocular depth $\hat{D}$ as a normal estimation $\mathbf{\hat{N}}$ to supervise rendered Gaussian normal $\mathbf{\bar{N}}$:
$\mathcal{L}_{normal}=\sum(1-\mathbf{\hat{N}}\cdot\mathbf{\bar{N}}).$ (11)
A Total Variation (TV) regularization [45] is applied to the rendered depth maps $\bar{D}$ to encourage surface smoothness:
$\mathcal{L}_{tv}=\sum_{i,j}g_{c}\cdot(|\bar{D}_{i+1,j}-\bar{D}_{i,j}|+|\bar{D}_{i,j+1}-\bar{D}_{i,j}|),$ (12)
where $i,j$ are the indices of pixels, $g_{c}=\exp(-\nabla\mathbf{C})$ is the gradient of RGB image. Therefore, our geometry regularization is:
$\mathcal{L}_{geo}=\lambda_{depth}\mathcal{L}_{depth}+\lambda_{normal}\mathcal{L}_{normal}+\lambda_{tv}\mathcal{L}_{tv}.$ (13)
The overall training loss is defined as a combination of scene representation loss and geometry regularization:
$\mathcal{L}=\mathcal{L}_{scene}+\mathcal{L}_{geo}.$ (14)
@@ 4.1 Experimental Settings
Baselines. We compare our proposal against several state-of-the-art methods. Among implicit methods, we compare against NeRF [18], VolSDF [48], NeuS [49], N-angelo [50], GeoNeus [51], Deep Blending [52], Instant NGP [53] and MipNeRF 360 [22]. As for explicit Gaussian methods, we compare against 3DGS [3], SuGaR [15], Mip-Splatting [23], 2DGS [1], GSurfel [54], 3DGSR [39] and GOF [2].
@@
Implementation. We perform single GPU training (NVIDIA 3090) and use by default $\lambda_{d}=100,\lambda_{n}=0.05,\lambda_{ns}=1000,\lambda_{fs}=10,\lambda_{depth}=0.1,\lambda_{normal}=0.1,\lambda_{tv}=1$. Regarding the multi-resolution hash grid settings, we set the number of hash-grid level as 20 and voxel size as 0.01. We set truncation during SDF training as 0.02. For more implementation details, please refer to supp. mat..
Mesh Extraction. Our method supports different methods for mesh extraction, e.g., TSDF fusion [14], Marching Cubes [56] and Marching Tetrahedral [16]. For evaluation, we adopt Marching Tetrahedral [16] approach for fast mesh extraction based on SDF. In contrast to GOF [2] which requires ray tracing to build the opacity field during extraction, ours does not require any additional step and can be performed within two minutes instead of hours.
@@ 4.2 Geometry Evaluation
We begin by comparing our method with state-of-the-art surface reconstruction techniques on the DTU dataset [46]. As shown in Table 1, we report the Chamfer Distance, where our approach outperforms all existing 3DGS-based and implicit methods, achieving superior reconstruction quality and efficiency. For N-angelo [50], we follow prior works and report the reproduced results from [57] as the published results cannot be reproduced. We visualize the qualitative results on DTU in Figure 3. Rendered normal maps from reconstructed meshes highlight that 2DGS [1] and GOF [2] often produce noisy, incomplete surfaces with numerous outliers and struggle with flat surface reconstruction. Our method addresses these issues by constraining Gaussians to the surface with SDF, resulting in smoother, more accurate reconstructions. It demonstrates that our method significantly improves surface reconstruction, particularly in challenging areas like flat surfaces or areas covered by sparse camera views.
We then compare our method on the Tanks and Temples dataset [47]. The evaluation is focused exclusively on foreground objects, as the ground truth point cloud does not cover background regions. As shown in Table 2, our method achieves results comparable to leading implicit approaches [50], while offering significantly faster optimization, training large unbounded scenes within 4 hours, compared to over 12 hours required by implicit methods. Additionally, our method achieves slightly improves results compared to GOF [2] with similar rendering efficiency and faster mesh extraction speed as shown in Table 4. Figure 4 provides a qualitative comparison, displaying rendered normal maps alongside ground truth color images. In the Barn scene, our method excels in reconstructing the flat rooftop where 2DGS [1] and GOF [2] struggle due to limited top-down camera views. The underlying continuity of the SDF field predicted by an MLP in our method enables accurate inference in this area. In the Caterpillar and Truck scenes, our approach successfully reconstructs reflective and transparent windows, a challenging task where both 2DGS [1] and GOF [2] fall short.
Figure 4: Surface Reconstruction on the Tanks and Temples Dataset [47]. We show the rendered normal map from reconstructed meshes. Resuls of 2DGS [1] and GOF [2] struggles to reconstruct smooth surfaces around transparent windows and produces noisy results in regions with sparse views. In contrast, our approach effectively handles transparent areas and achieves smooth reconstruction even in sparsely viewed regions.
@@
Figure 5: Surface Reconstruction on the Mip-NeRF 360 Dataset [22]. We show the rendered normal map from reconstructed meshes. GOF’s [2] mesh is noisy and it fails to reconstruct smooth surface where there are reflection and curvature surface. In contrast, our method can generate smoother and detailed surface without any outlier points.
@@ 4.3 Novel View Synthesis
We further compare our method with state-of-the-art novel view synthesis (NVS) methods on the Mip-NeRF 360 dataset [22]. Quantitative results in Table 3 indicate that our approach surpasses 2DGS [1] across all metrics in indoor scenes, while for outdoor scenes, the performance gap remains relatively small. However, our method significantly improves surface reconstruction quality. This suggests that our enhanced geometry representation leads to higher visual quality. Figure 5 presents a qualitative comparison of the extracted meshes. Consistent with our observations on the Tanks and Temples dataset [47], our method reconstructs smooth, detailed meshes. In contrast, 2DGS [1] yields overly smooth meshes, and GOF [2] produces noisy artifacts on flat surfaces.
@@ 4.4 Ablation Study
Figure 6: Ablation Study.
4.4.1 Mesh Extraction Comparison
Our method supports mesh extraction from 3D Gaussians using a variety of techniques. All mesh extractions are performed on a single NVIDIA 3090 GPU with 24GB of memory. To apply the Marching Cubes algorithm in unbounded scenes, we follow the 2DGS approach from [1], compressing the unbounded volume into a bounded range. However, the Marching Cubes algorithm is memory-intensive; on a 24GB GPU, it supports only low-resolution reconstructions, limiting its feasibility in unbounded scenes. For visualization purposes, we crop the Barn scene from the Tanks and Temples dataset [47] to a smaller range before running Marching Cubes. As shown in Figure 6, our method is still able to generate fine-detailed meshes with Marching Cubes.
In addition to Marching Cubes, we use Marching Tetrahedra for mesh extraction, which is not limited by resolution. In Table 4, we report the mesh extraction time of different methods, demonstrating the effectiveness of our Marching Tetrahedra approach. 2DGS uses TSDF fusion, reconstructing meshes in 10-30 minutes depending on resolution, and GOF requires ray tracing to distill opacity fields from the 3DGS model before applying their own version of Marching Tetrahedra. In contrast, our approach leverages the SDF and 3DGS to initialize tetrahedral points, achieving high-detail mesh reconstruction within 2 minutes.
Meshing Time (min) TSDF M.Cube M.Tetrahedral
2DGS $\sim$10-30 × ×
GOF $\sim$10-30 × $\sim$30
Ours $\sim$10-30 $\sim$10-30 $\sim$2
F1 Score w/o Geo. w/ Geo.
Ours 0.56 0.62
Table 4: Quantitative Ablation Study. We show the mesh extraction methods and time. We denote Marching Cube [56] as M.Cube and Marching Tetrahedral [16] as M.Tetrahedral. We also report the F1 score of Truck reconstruction with and without proposed geometry regularization.
4.4.2 Geometry Regularization
We perform an ablation study of the proposed geometry regularization in the Truck scene from the Tanks and Temples dataset [47] which has transparent windows that are normally hard to reconstruct. Figure 6 shows visual comparisons with and without the use of a depth estimator as a pseudo-ground-truth depth map, highlighting that pseudo geometry regularization significantly improves the reconstruction of flat and transparent surfaces. Quantitative results in Table 4 further demonstrate the effectiveness of this geometry regularization technique.
@@ 5 Conclusion
We present G2SDF, a novel method for efficient, high-quality surface reconstruction that combines explicit Gaussian representations with implicit SDFs. By normalizing unbounded coordinates into a fixed range, we enable efficient positional embedding. Using a hash-grid-based SDF, we predict SDF values for each Gaussian and convert them to opacity values for rasterization, effectively aligning Gaussians to the surface for accurate geometry representation. G2SDF enables fast and detailed mesh extraction. Experimental results show that G2SDF matches state-of-the-art methods in both surface reconstruction and novel view synthesis, producing high-quality meshes with fine details.
@@ A.1 Hyperparameters
Gaussian Splatting. For hyperparameters used in the Gaussian Rasterization, we follow previsous works [2, 1, 3]. We train our Gaussian model and SDF network jointly with 30000 iterations. We set the initial learning rate for Gaussians’ position as 0.00016 and the final initial learning rate as 0.0000016. And we set learning rates for opacity, scales and rotation as 0.05, 0.005, 0.001, respectively. We start the Gaussian densification after 500 iterations and until 15000 iterations. We densify Gaussians every 100 iterations and the gradient threshold for densification is 0.0002. We start the distortion loss $\mathcal{L}_{d}$ and the depth-normal loss $\mathcal{L}_{n}$ after 3000 iterations and 7000 iterations, respectively.
SDF Network. We use two-layer MLPs and the hidden dimension is 32. We initialize $\beta$ from Eq. 6 as 100 and set learning rate as 0.005. We start train our SDF from 5000 iterations and until 30000 iterations. The learning rate for our SDF network is 0.002. For each iteration, we sample $M=10000$ pixels and $K_{n}=11,K_{f}=64$ points.
Geometry Regularization. We set the learning rate for $a,b$ in Eq. 10 as 0.1. and for Tanks and Temples, we set $\lambda_{depth}=0.1,\lambda_{normal}=0.1$, for DTU and Mip-NeRF 360, we set $\lambda_{depth}=0.01,\lambda_{normal}=0.01$.

### v2

@@ 3.3 Optimization
More specifically, we first sample $M$ pixels $\{\mathbf{u}=\mathbf{u}_{i}\ |\ \mathbf{u}_{i}\in\mathbb{R}^{2},{\text{for}\ i=1,2,...,M\}}$ from the rendered Gaussian depth map $\bar{D}$. For each ray casts from a pixel, we sample $K=K_{n}+K_{f}$ points $\{t_{k}\}$ from the camera center to the surface, where $K_{n}$ means near surface samples and $K_{f}$ means free space samples. SDF values of each point can be queried as $\{\bar{s}_{k}\}$. For near surface samples within the truncation region $tr$, i.e. $S_{tr}=\{|\bar{D}(\mathbf{u})-t_{k}|\leq tr\}$, we use the distance between the sampled point $t_{k}$ and its surface $\bar{D}(\mathbf{u}_{j})$ as an approximation of ground truth SDF value for supervision:
@@
For points that are far from the surface $S_{fs}=\{\bar{D}(\mathbf{u})-t_{k}>tr\}$, we apply a free-space supervision as:
@@
Therefore, the overall SDF regularization is $\lambda_{sdf}=\lambda_{ns}\mathcal{L}_{ns}+\lambda_{fs}\mathcal{L}_{fs}$.
Multi-Resolution Regularization. Unlike vanilla 3DGS [16] that uses L1 and SSIM losses with opacity-based pruning, our method links opacity to spatial position, lifting the need for Opacity Resets but potentially causing Gaussian redundancy. While low-to-high frequency strategies [52] can accelerate convergence, they risk 3D structural blurring. Instead, we propose multi-resolution supervision through wavelet transforms, progressively increasing resolution while preserving structural integrity:
${\mathbf{C}_{l}}=\mathcal{W}({\mathbf{C}},l)$ (11)
where $\mathbf{C}$ is ground truth color images, $\mathcal{W}$ represents wavelet transform and $l$ denotes the level of transformation. This approach avoids the zigzag artifacts of direct image rescaling, ensuring stable detail refinement during training. It enables the Gaussian model to first capture a rough scene representation with fewer Gaussians and later refine scene details for a more precise Gaussian distribution. Therefore, the photometric regularization is $\mathcal{L}_{p}=0.8*|{\mathbf{C}_{l}}-\bar{\mathbf{C}}|+0.2*SSIM({\mathbf{C}_{l}},\bar{\mathbf{C}})$.
Objective Function. During training, we jointly optimize explicit Gaussians and implicit SDF. We apply distortion $\mathcal{L}_{d}$ and depth-normal $\mathcal{L}_{n}$ regularization as defined in GOF [48]. The overall objective function is defined as:
$\mathcal{L}=\mathcal{L}_{p}+\mathcal{L}_{geo}+\mathcal{L}_{sdf}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{n}.$ (12)
Pruning. In our framework, where opacity is intrinsically linked to spatial position through the SDF relationship, we implement a geometry-aware pruning strategy that removes Gaussians exceeding the truncation threshold: $\bar{s}_{i}>tr$. This approach replaces conventional opacity-based pruning with a physically-grounded criterion that better aligns with our surface-aware representation.
Mesh Extraction. Previous Marching Cubes-based approaches [45, 24] must densely search both occupied and free space, leading to inefficiencies and resolution limitations, especially in unbounded scenes with coordinate distortions. In contrast, our method adapts Marching Tetrahedral [18] and exploits 3DGS’s spatial cues to constrain the search to regions only near Gaussians, avoiding unnecessary exploration of empty space while achieving superior surface alignment and reconstruction efficiency.
@@ 4.1 Experimental Settings
Baselines. We compare our proposal against several state-of-the-art methods. Among implicit methods, we compare against NeRF [25], VolSDF [44], NeuS [37], N-angelo [21], GeoNeus [10], Instant NGP [27] and MipNeRF 360 [3]. As for explicit Gaussian methods, we compare against 3DGS [16], SuGaR [11], Mip-Splatting [47], 2DGS [12], GSurfel [8], 3DGSR [24], GS-Pull [55] and GOF [48].
@@
Implementation. We perform single GPU training (NVIDIA 3090) and use by default $\lambda_{d}=100,\lambda_{n}=0.05,\lambda_{ns}=1000,\lambda_{fs}=10,\lambda_{depth}=0.05,\lambda_{normal}=0.1$ and $\beta=100$. We train 30000 iterations like other methods and do not require extra training for SDF. For more implementation details, please refer to supp. mat..
@@
Figure 3: Surface Reconstruction on the DTU Dataset [1]. We show normal maps rendered from the reconstructed meshes. 2DGS [12] produces overly smooth surfaces with floating artifacts and outliers, while GOF [48] struggles to reconstruct flat surface. In contrast, our method generates smooth yet detailed surfaces without floating artifacts, demonstrating its superior ability to represent geometry accurately.
@@ 4.2 Geometry Evaluation
We evaluate our method on the DTU dataset [13], comparing it with state-of-the-art surface reconstruction techniques. As shown in Table 1, our approach achieves the lowest Chamfer Distance, surpassing all existing 3DGS-based and implicit methods in both quality and efficiency. For N-angelo [21], we report results reproduced in [35], as the original values could not be verified. Figure 3 presents qualitative comparisons, where normal maps reveal that 2DGS [12] and GOF [48] often produce noisy, incomplete surfaces with artifacts, particularly struggling with flat regions. Our method mitigates these issues by constraining Gaussians to the surface via SDF, yielding smoother and more accurate reconstructions, especially in challenging areas with sparse camera coverage.
@@
We then compare our method on the Tanks and Temples dataset [17]. The evaluation is focused exclusively on foreground objects, as the ground truth point cloud does not cover background regions. As shown in Table 2, our method achieves results on par with leading implicit approaches [21] while significantly reducing training time—optimizing large unbounded scenes in around 3 hours versus over 12 hours for implicit methods. Additionally, we slightly outperform GOF [48] while maintaining comparable rendering efficiency and faster mesh extraction (Table 4). Figure 4 showcases qualitative results, where our method excels at reconstructing flat surfaces, such as the Barn rooftop, and handling reflective and transparent windows in the Caterpillar and Truck scenes—areas where 2DGS [12] and GOF [48] struggle due to limited views and scene complexity. Our Gaussian-guided implicit surface reconstruction method ensures smooth and accurate reconstructions in these challenging scenarios.
Figure 4: Surface Reconstruction on the Tanks and Temples Dataset [17]. We show the rendered normal maps from reconstructed meshes. Resuls of 2DGS [12] and GOF [48] struggles to reconstruct smooth surfaces around transparent windows and produces noisy results in regions with sparse views. In contrast, our approach effectively handles transparent areas and achieves smooth reconstruction even in sparsely viewed regions.
Figure 5: Surface Reconstruction on the Mip-NeRF 360 Dataset [3]. We show the rendered normal maps from reconstructed meshes. GOF’s [48] meshes are noisy and fail to reconstruct smooth surface where there are reflection and curvature surface. In contrast, our method can generate smoother and detailed surface without any outlier points.
@@ 4.3 Novel View Synthesis
We further compare our method with state-of-the-art novel view synthesis (NVS) methods on the Mip-NeRF 360 dataset [3]. Quantitative results in Table 3 indicate that our approach surpasses GOF [48] across all metrics in indoor scenes, while for outdoor scenes, the performance gap remains relatively small. However, our method significantly improves surface reconstruction quality. This suggests that our enhanced geometry representation leads to higher visual quality. Figure 5 presents a qualitative comparison of the extracted meshes. Consistent with our observations on the Tanks and Temples dataset [17], our method reconstructs smooth, detailed meshes. In contrast, 2DGS [12] yields overly smooth meshes, and GOF [48] produces noisy artifacts on flat surfaces.
@@ 4.4 Ablation Study
Figure 6: Ablation Study on Geometry Regularization.
4.4.1 Runtime Analysis
Table 4 presents a comprehensive runtime analysis on the Barn scene, highlighting our method’s efficiency and adaptability. Notably, our framework seamlessly integrates with existing Gaussian rasterizers, as demonstrated by the “Ours (2D)” implementation combining our approach with the 2DGS rasterizer. While the incorporation of SDF guidance enables superior surface reconstruction quality, our method maintains competitive training times comparable to efficient baselines like 2DGS and GOF. Furthermore, our approach achieves substantial acceleration in mesh extraction compared to existing methods.
4.4.2 Ablations on Different Regularizations
Ablations on SDF Fields. Table 5 quantitatively validates the effectiveness of our Gaussian-based implicit surface reconstruction framework compare with GOF. To further demonstrate the generalizability of our approach, we additionally evaluate its performance when integrated with the 2DGS rasterizer denoted as “Ours (2D)”, reporting consistent F1-score improvements across different configurations.
Run Time 2DGS Ours (2D) GOF Ours
Training 17 min 39 min 2.8 h 3.1 h
Meshing $\sim$20 min 5 min $\sim$30 min 5 min
Table 4: Runtime. We report the training and meshing time.
Method 2DGS Ours (2D) $\lambda_{D}=0$ $\lambda_{N}=0$ $\lambda_{D}=0$, $\lambda_{N}=0$
F1 Score 0.36 0.40 0.52 0.51 0.50
Method GOF GOF+Geo $\lambda_{D}=0.5$ $\lambda_{N}=0.5$ Ours
F1 Score 0.51 0.52 0.45 0.47 0.56
Table 5: Quantitative Ablation Study. We report F1 score of scene Barn with and without proposed regularization. We denote our method with default 3D Gaussian rasterizer [48] as “Ours” and with 2D Gaussian rasterizer [12] as “Ours (2D)”. “GOF+Geo” represents GOF with geometry cues only.
Ablations on Geometry Regularization. The ablation studies in Table 5 demonstrate the impact of different geometry regularization terms. Since the pseudo normal map is derived from the pseudo depth map, they exhibit identical performance gaps. We further evaluate GOF enhanced with our geometry cues, denoted as “GOF+Geo”, which shows modest improvements (0.52). Due to the scale ambiguity problem, we set small values for $\lambda_{D},\lambda_{N}$ specifically (0.05 and 0.1 for Tanks and Temples, 0.01 and 0.01 for DTU and MipNeRF 360). We demonstrate that increasing these weights to 0.5 actually degrades the results by introducing adverse effects from geometric cues. We also demonstrate the error distribution after alignment in supp. mat., showcasing the limitation of geometry cues. The most significant enhancement is achieved through our complete pipeline (8% improvement), highlighting the synergistic effect of our integrated approach. Figure 6 demonstrates the effectiveness of our geometry regularization in handling transparent and flat regions, significantly improving reconstruction quality through enhanced surface consistency.
@@ 5 Conclusion
We introduce MonoGSDF, a Gaussian-based implicit surface reconstruction framework. Our SDF-to-Opacity function bridges implicit and explicit representations, while Gaussian-guided normalization ensures stable optimization across unbounded 3D scenes. A multi-resolution training strategy, enhanced by the geometric cues, progressively refines reconstruction quality. Our method achieves state-of-the-art accuracy and efficiency, effectively handling complex cases such as flat and transparent surfaces where existing approaches often struggle.
@@ A.1 Hyperparameters
Gaussian Splatting. For hyperparameters used in the Gaussian rasterization, we follow previsous works [48, 12, 16]. We train our Gaussian model and SDF network jointly with 30000 iterations. We set the initial learning rate for Gaussians’ position as 0.00016 and the final initial learning rate as 0.0000016. And we set learning rates for scales and rotation as 0.005, 0.001, respectively. We start the Gaussian densification after 500 iterations and until 15000 iterations. We densify Gaussians every 100 iterations and the gradient threshold for densification is 0.0002. We start the distortion loss $\mathcal{L}_{d}$ and the depth-normal loss $\mathcal{L}_{n}$ after 3000 iterations and 7000 iterations, respectively. And we set the wavelet level as 3 and gradually increase the resolution until 10000 iteration. After that, we use the full resolution for training.
SDF Network. We use two-layer MLPs and the hidden dimension is 32. We initialize $\beta$ from Eq. 5 as 100. We start train our SDF from 5000 iterations and until 30000 iterations. The learning rate for our SDF network is 0.002. For each iteration, we sample $M=10000$ pixels and $K_{n}=11,K_{f}=64$ points.
Geometry Regularization. We set the learning rate for $s,t,a,b$ in Eq. 7 as 0.01. And for Tanks and Temples, we set $\lambda_{D}=0.05,\lambda_{N}=0.1$, for DTU and Mip-NeRF 360, we set $\lambda_{D}=0.01,\lambda_{N}=0.01$.